if F(x)=4-x and g(x)=2x+3, find value of x where F(x-2)=g(2x).
Answer by Nate(3500)

f(x - 2) = g(2x)
4 - (x - 2) = 2(2x) + 3
4 - x + 2 = 4x + 3
3 = 5x
3/5 = x